# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### 29 Nov 2023 v1

@@ 5 Conclusion
We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations. Our method, Thera, implicitly describes an image as a combination of sinusoidal components which can be modulated based on their frequency to perform Gaussian filtering (anti-aliasing) analytically with little overhead. Our experimental evaluation shows that Thera is superior to competing methods in most cases, while also being more parameter efficient and providing theoretical guarantees. We also hope to inspire others to apply Thera to other problems where it is likely to be beneficial due to its Gaussian observation model, such as demosaicing, burst super-resolution, and neural radiance fields.
Supplementary Material
Appendix A Theory
A.1 Preliminaries
As was described in Sec. 3.1, the idea is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
@@ A.1 Preliminaries
The reason for this is that the analytical solution to the (isotropic) heat equation can be modelled as a convolution of the initial state $\Phi(\mathbf{x},0)$ with a Gaussian kernel
@@
By fitting the data at $\Phi(\mathbf{x},1)$, we are assuming a Gaussian point spread function (PSF) with the shape
@@
The ill-posed problem for $10<t<$ 1 is the interesting case. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
@@
A.2 Thermal diffusivity coefficient
To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$. One way to do so is to match the cut-off frequency of the filter in Eq. 9 at $t=1$ to the well-known Nyquist frequency given by the sampling rate. We take the cut-off frequency of the Gaussian filter defined in Eq. 9 to be the frequency whose amplitude is halved, which is
@@ A.2 Thermal diffusivity coefficient
where $N$ is the number of samples along a given dimension. This formulation assumes even sampling over $x_{1}$ and $x_{2}$.
@@
For the Thera formulation presented in Sec. 3.2, we want $\Phi$ to contain a single pixel at $t=1$, therefore we initialize $\kappa$ with
@@
Note that the exact value of $\kappa$ will depend on the exact characteristics of the system that is being modelled. Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but is also more prone to aliasing.
Finally, we would like to highlight that Eq. 12 is specific to the case where $\mathbf{x}$ is 2-dimensional. The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using fields in spaces with numbers of spatial dimensions other than 2. Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the desired formulas for a Gaussian filter with the appropriate number of dimensions.
A.3 Relationship between $t$ and $s$
Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate. If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it with $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$. Using Eq. 12 and Eq. 13, we can easily find the quadratic relationship
@@ A.3 Relationship between $t$ and $s$
Thus, $10<t<$ 1 refer to super-resolution, while $t>1$ refers to downsampling.
@@
A.4 Other filters
In theory, the formulation presented in Sec. 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely. Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation. Initial explorations of a sharp low-pass filter which completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely to how this also affected gradients during training. It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way. For quantitative evaluations this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios this may be desirable.
A.5 Initialization of components
We have noticed during our experiments that the initialization of the components ($\mathbf{W}_{1}$ in Eq. 2) is very important. Sitzmann et al. [38] made similar observations when periodic activation functions were first used for neural fields. The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training. Thus, we chose to initialize $\mathbf{W}_{1}$ such that
Appendix B Additional qualitative results
Figure 8 and Fig. 9 contain additional qualitative results for lower and higher scaling factors, respectively. Again, we see that Thera can reconstruct the targets more faithfully. This is particularly visible in repetitive structural elements. We suspect that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this.
Appendix C Additional quantitative results
Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab. 1. We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric. Furthermore, Tab. 6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$). Finally, in Fig. 7 we show the number of additional parameters for various methods on the ×$\times 3$ upsampling task.
Appendix D Hypernetwork architecture
@@
In Fig. 6, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L. The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image. It is then followed by a number of (residual) ConvNeXt [25] blocks. Each block consists of i) a $3\times 3$ convolution that maps to $d$-dimensional features, ii) a layer normalization [2] operation, iii) a $1\times 1$ convolution with $4d$ filters, iv) a GELU [15] non-linearity, and v) another $1\times 1$ convolution with $d$ filters. Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al. [25] – to $3\times 3$ to reduce the number of parameters. Furthermore, note that in Thera-L, the first convolution per block uses $d$ groups (effectively implementing a channel-wise convolution), while for the two smaller variants we used full convolutions, i.e., a single group. In our implementation, Thera-S uses zero ConvNeXt blocks, Thera-M uses 5 blocks with $d=64$, and Thera-L uses 6 blocks with $d=64$ followed by 7 blocks with $d=96$ and 3 blocks with $d=128$. For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
@@
Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [38]. There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L. Let $c$ be the number of components used (32, 512 and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork. The total parameter count of the hypernetwork is 8,192, 0.48 M and 1.41 M for the three variants, respectively. To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters. We highlight that these do not come from the hypernetwork.
Appendix E Replication of other methods
@@
We encountered challenges attempting to recreate results of some of the competitor methods, which explains why some numbers are missing in various tables or differ from the originally reported ones. Details are provided below.
@@
CUF [41]. At the time of writing, there was no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples and could not include CUF in the multi-scale consistency study in Sec. 4.2.
@@
CLIT [10]. For CLIT, at the time of writing, there is public code available, however, checkpoints are not. We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU hours: over 1 month using 8×8$\times$8 × Nvidia GeForce RTX 3090s. The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
@@
CiaoSR [8]. We noticed that CiaoSR’s border cropping prior to evaluation on the DIV2K dataset in the official implementation deviated slightly from the rest of the comparison methods. We aligned this to align the evaluation of all methods and re-computed all DIV2K numbers. This explains the slight deviation from the original paper.
Figure 7: Plots depicting the number of additional parameters vs. PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 3$ upsampling task.
@@
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
@@
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth

### 14 Mar 2024 v2

@@ 5 Conclusion
We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations. Our method, Thera, implicitly describes an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) analytically, with negligible overhead. Our experimental evaluation shows that Thera outperforms competing methods in most cases, while at the same time being more parameter efficient, and offering theoretical guarantees. We are convinced that Thera-style representations could also benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful observation models like our Gaussian PSFs.
Appendix 0.A Theory
0.A.1 Preliminaries
As was described in Sec. 3.1, the idea underlying our neural field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
@@ 0.A.1 Preliminaries
The reason for this is that the analytical solution to the (isotropic) heat equation can be modeled as a convolution of the initial state $\Phi(\mathbf{x},0)$ with a Gaussian kernel
@@
By fitting the data (image $I$) at $\Phi(\mathbf{x},1)$, we are assuming a Gaussian point spread function (PSF) with the shape
@@
The ill-posed problem for $10<t<$ 1 is the interesting case, where this formulation relates to super-resolution. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
@@
0.A.2 Thermal Diffusivity Coefficient
To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$. One way to do so is to match the cut-off frequency of the filter in Eq. 9 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate. We take the cut-off frequency of the Gaussian filter defined in Eq. 9 to be the frequency whose amplitude is halved, which is
@@ 0.A.2 Thermal Diffusivity Coefficient
where $N$ is the number of samples along a given dimension. This formulation assumes even sampling over $x_{1}$ and $x_{2}$. To extend this formulation to non-square images, it would be necessary to change the shape of the signal’s domain in order to maintain even sampling in all spatial dimensions.
@@
For the Thera formulation presented in Sec. 3.2, we want $\Phi$ to contain a single pixel at $t=1$. This is the pixel from the low-resolution input which will become $SR^{2}$ pixels for super-resolution with a scaling factor of $SR$. Therefore we initialize $\kappa$ with
@@
Note that the exact value of $\kappa$ will depend on the characteristics of the system that is being modeled and the anti-aliasing filter that was used (or is assumed). Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but are also more prone to aliasing.
Finally, we would like to highlight that Eq. 12 is specific to the case where $\mathbf{x}$ is 2-dimensional. The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using our field’s formulation in spaces with numbers of spatial dimensions other than 2. Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the formulas for a Gaussian filter with the appropriate number of dimensions.
0.A.3 Relationship Between $t$ and $s$
Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate. If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it down to $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$. Using Eq. 12 and Eq. 13, we can easily find the quadratic relationship
@@ 0.A.3 Relationship Between $t$ and $s$
Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling. This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
0.A.4 Other Filters
In theory, the formulation presented in Sec. 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely. Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation. Initial explorations of a sharp low-pass filter that completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely due to the associated effect on gradients during training. It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way. For quantitative evaluations, this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios, this may be desirable.
0.A.5 Initialization of Components
We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Eq. 2, is important. Sitzmann et al. [40] made similar observations when periodic activation functions were first used for neural fields. The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training. Thus, we choose to initialize $\mathbf{W}_{1}$ such that
Appendix 0.B Additional Qualitative Results
Figures 9 and 10 contain additional qualitative results for lower and higher scaling factors, respectively. Again, we see that Thera can reconstruct the targets more faithfully. This is particularly (but not exclusively) visible in repetitive structural elements. It seems evident that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this. Furthermore, we provide a showcase of the continuous upsampling capabilities of our method in Fig. 6.
Figure 6: Showcase of multiscale upsampling using Thera-L with a SwinIR [23] backbone, shown with non-integer scaling factors.
Appendix 0.C Additional Quantitative Results
Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab. 1. We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric. Furthermore, Tab. 6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$). Finally, in Fig. 7 we compare the number of additional parameters and PSNR values for various methods and upsampling factors on DIV2K.
Figure 7: Plots depicting the number of additional parameters vs. PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 2$ (left), ×$\times 3$ (center) and ×$\times 4$ (right) upsampling tasks.
@@ Appendix 0.C Additional Quantitative Results
Appendix 0.D Hypernetwork Architecture
In Fig. 8, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L. The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image. It is then followed by a number of (residual) ConvNeXt [26] blocks. Each block consists of (i) a $3\times 3$ convolution that maps to $d$-dimensional features, (ii) a layer normalization [2] operation, (iii) a $1\times 1$ convolution with $4d$ filters, (iv) a GELU [15] non-linearity, and (v) another $1\times 1$ convolution with $d$ filters. Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al. [26] – to $3\times 3$ to reduce the number of parameters. Furthermore, note that in Thera-L, the first convolution per block uses $d$ groups (effectively implementing a channel-wise convolution), while for the two smaller variants we used full convolutions, i.e., a single group. In our implementation, Thera-S uses zero ConvNeXt blocks, Thera-M uses 5 blocks with $d=64$, and Thera-L uses 6 blocks with $d=64$ followed by 7 blocks with $d=96$ and 3 blocks with $d=128$. For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
@@ Appendix 0.D Hypernetwork Architecture
Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [40]. There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L. Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork. The total parameter count of the hypernetwork is 8,192, 0.48 M, and 1.41 M for the three variants, respectively. To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters. We highlight that these do not come from the hypernetwork.
Appendix 0.E Replication of Other Methods
@@
We encountered challenges attempting to recreate results of some of the competing methods, which explains why some numbers are missing in various tables or differ from the originally reported ones. Details are provided below.
@@
CUF [43]. At the time of writing, there were no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples and could not include CUF in the multi-scale consistency study in Sec. 4.2.
@@
CLIT [10]. For CLIT, at the time of writing, there is public code available, however, checkpoints are not. We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU time: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090s. The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
@@
CiaoSR [8]. We noticed that in the official CiaoSR implementation, border cropping prior to evaluation on the DIV2K dataset deviated slightly from all other methods. We aligned this to enable a meaningful comparison and re-computed all DIV2K numbers. This explains the slight deviation from the original paper.
@@
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
@@
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
@@
Appendix 0.F Computational Requirements
For comparison of computational requirements between methods, we provide training time, inference time, and inference VRAM requirements in Tab. 7, for the EDSR-baseline backbone. These values assume NVIDIA GeForce RTX 3090 Ti GPUs, and they show that we are more efficient than recent transformer-based methods, but still reach superior performance, as shown in the main paper.
Table 7: Computational requirements of various methods.
w/ EDSR-b Training time Inference time VRAM
(GPU days)
MetaSR 0.42 3.68 ms 588 MiB
LIIF 0.71 13.7 ms 1174 MiB
LTE 0.79 17.1 ms 1250 MiB
CiaoSR $\sim 2.6$ 37.7 ms 6474 MiB
CLIT $\sim 4.6$ 16.0 ms 2984 MiB
Thera-S 1.0 3.52 ms 754 MiB
Thera-M 2.2 10.0 ms 2290 MiB
Thera-L 3.3 14.3 ms 2292 MiB